# Bath-induced interactions and transient dynamics in open quantum systems at strong coupling: Effective Hamiltonian approach

### v1

@@ IV Assessment of the effective Hamiltonian method
Given its underlying approximations, the effective Hamiltonian method is useful when the temperature of the bath is comparable to energy levels in the system, while the characteristic frequency of the bath is high, i.e., $\Omega\gg\Delta,T$. However, as was demonstrated in Ref. Anto-Sztrikacs et al. (2023b), performing the polaron mapping in Eq. (4) in a variational manner can extend the reach of the method to handle cases with $\Omega$ comparable to $\Delta$ and $T$.
V Conclusions
@@
In delimiting the scope of the problem, we have adopted specifically tailored initial states and a bath spectral density function. Our working assumption is the one required by the applicability of a Markovian embedding via the reaction-coordinate mapping Iles-Smith et al. (2014). This assumption requires specific spectral density functions such that after the mapping the residual reservoir is coupled weakly to the newly-mapped Hamiltonian. Aside from that, the assumption on the energy scale imposed by the bath characteristic energy $\Omega\gtrsim\lambda,T$ remains valid in our calculations. Interestingly, such condition may be relaxed towards understanding equilibrium states as was recently demonstrated in Refs. Anto-Sztrikacs et al. (2023a, b); Min et al. (2023). However, for transient dynamics, relaxing this assumption will introduce more complicated transitions between the system and bath that further complicate the dynamics.
@@
We have found that the coupling of a two-spin system to a common thermal reservoir leads to correlations between each spin even if the spins are not coupled to each other. Interestingly, at equilibrium these are all classical correlations but the transient states develop non-classical correlations stemming from non-separability. As a function of temperature, non-classical correlations between each spin subsystem may develop at low temperature in the steady state regime. We leave an investigation of the temperature dependence of correlations to future work.
Our results may be extended by considering a less extreme truncation scheme for the reaction coordinate manifold. Instead of retaining only the ground state of the reaction coordinate, one may keep two or more levels in the effective Hamiltonian. This extended treatment is expected to improve on the predicted dynamics. However, it will complicate the procedure and its analysis, requiring working with a larger space and tracing out the reaction coordinate as part of the procedure.
@@ Appendix C Dynamics from the hierarchical equations of motion
In general, in arriving at the Redfield equation Eq. (II.3) one relies on the Born-Markov approximation, which does not guarantee that the resulting equation leads to an appropriate definition of a generator of a dynamical semigroup. In turn, this may translate to non-positive dynamics of the reduced density matrix of the state. Though not typically an issue at weak system-reservoir coupling, non-positive dynamics may pose an issue in our calculations. In this Section, we however demonstrate that the dynamics generated by Eq. (II.3) is correct when compared to another numerically exact approach: the hierarchical equations of motion. We remark that our Markovian embedding introduced in Sec. II.2 guarantees that the enlarged system remains weakly coupled to the residual reservoir and, therefore, one would not expect positivity of the density matrix to be violated along the time trajectory generated by Eq. (II.3). However, we find it important to confirm whether the correct dynamics are being generated in our model.
@@
Though, in general, the Redfield equation does not preserve complete positivity of the reduced density operator along time trajectories, we find that in our case the true dynamics are indeed captured for our particular system and bath configuration.

### v2

@@ V Assessment of the effective Hamiltonian method
Given its underlying approximations, the effective Hamiltonian method is useful when the temperature of the bath is comparable to energy levels in the system, while the characteristic frequency of the bath is high, i.e., $\Omega\gtrsim\Delta,T$. However, as was demonstrated in Ref. Anto-Sztrikacs et al. (2023b), performing the polaron mapping in Eq. (4) in a variational manner can extend the reach of the method to handle cases with $\Omega$ comparable to $\Delta$ and $T$.
VI Conclusions
@@
In delimiting the scope of the problem, we have adopted specifically tailored initial states and a bath spectral density function. Our working assumption is the one required by the applicability of a Markovian embedding via the reaction-coordinate mapping Iles-Smith et al. (2014). This assumption requires specific spectral density functions such that after the mapping the residual reservoir is coupled weakly to the newly-mapped Hamiltonian. Aside from that, the assumption on the energy scale imposed by the bath characteristic energy $\Omega\gtrsim\lambda,T$ remains valid in our calculations. Interestingly, such condition may be relaxed towards understanding equilibrium states as was recently demonstrated in Refs. Anto-Sztrikacs et al. (2023a, b); Min et al. (2023).
@@
We have found that the coupling of a two-spin system to a common thermal reservoir leads to correlations between each spin even if the spins are not directly coupled to each other. Interestingly, at equilibrium these are all classical correlations but the transient states develop non-classical correlations stemming from non-separability. As a function of temperature, non-classical correlations between each spin subsystem may develop at low temperature in the steady state regime. We leave an investigation of the temperature dependence of correlations to future work.
Our results may be extended by considering a less extreme truncation scheme for the reaction coordinate manifold. Instead of retaining only the ground state of the reaction coordinate, one may keep two or more levels in the effective Hamiltonian. This extended treatment is expected to improve on the predicted dynamics. However, it will complicate the procedure and its analysis, requiring working with a larger space and tracing out the reaction coordinate as part of the procedure. More broadly, it is interesting to explore alternative mapping approaches complementary to the effective Hamiltonian: methods that are based on the Schriffer-Wolff transformation, chain mapping, or alternative decomposition of bath modes, see e.g. Refs. Prior et al. (2010); Chin et al. (2010); Woods et al. (2014); Ashida et al. (2021, 2022); Cirio et al. (2023).
@@ Appendix C Dynamics from the hierarchical equations of motion
In general, in arriving at the Redfield equation Eq. (II.3) one relies on the Born-Markov approximation, which does not guarantee that the resulting equation leads to an appropriate definition of a generator of a dynamical semigroup. In turn, this may translate to non-positive dynamics of the reduced density matrix of the state. Though not typically an issue at weak system-reservoir coupling, non-positive dynamics may pose an issue in our calculations. Furthermore, the Redfield equation neglects non-Markovian effects, which may be important if the residual coupling is not weak.
In this Section, we however demonstrate that the dynamics generated by Eq. (II.3) is correct when compared to another numerically exact approach: the hierarchical equations of motion. We remark that our Markovian embedding introduced in Sec. II.2 guarantees that the enlarged system remains weakly coupled to the residual reservoir and, therefore, one would not expect positivity of the density matrix to be violated along the time trajectory generated by Eq. (II.3). However, we find it important to confirm whether the correct dynamics are being generated in our model.
@@
Though, in general, the Redfield equation does not preserve complete positivity of the reduced density operator along time trajectories, and it ignores non-Markovian effects, we find that in our case the true dynamics are indeed captured for the reaction coordinate Hamiltonian, for our particular system and bath configuration.